### v2

# Integrating Multi-Modal Input Token Mixer into Mamba-Based Decision Models: Decision MetaMamba

Abstract
Sequence modeling with State Space models (SSMs) has demonstrated performance surpassing that of Transformers in various tasks, raising expectations for their potential to outperform the Decision Transformer and its enhanced variants in offline reinforcement learning (RL). However, decision models based on Mamba, a state-of-the-art SSM, failed to achieve superior performance compared to these enhanced Decision Transformers. We hypothesize that this limitation arises from information loss during the selective scanning phase. To address this, we propose the Decision MetaMamba (DMM), which augments Mamba with a token mixer in its input layer. This mixer explicitly accounts for the multimodal nature of offline RL inputs, comprising state, action, and return-to-go. The DMM demonstrates improved performance while significantly reducing parameter count compared to prior models. Notably, similar performance gains were achieved using a simple linear token mixer, emphasizing the importance of preserving information from proximate time steps rather than the specific design of the token mixer itself. This novel modification to Mamba’s input layer represents a departure from conventional timestamp-based encoding approaches used in Transformers. By enhancing performance of Mamba in offline RL, characterized by memory efficiency and fast inference, this work opens new avenues for its broader application in future RL research.
Codehttps://github.com/too-z/decision-metamamba
Introduction
Figure 1: the number of parameters per score for each dataset
Mamba(Gu and Dao 2023) is recognized as one of the most advanced state-space models (SSMs), demonstrating superior performance over Transformers in various sequence modeling domains such as language processing, vision(Zhu et al. 2024)(Liu et al. 2024), graph analysis(Wang et al. 2024a), and time series(Wang et al. 2024c)(Patro and Agneeswaran 2024b). Mamba’s architecture allows its internal states to change input-dependently, enabling selective information retention and facilitating content-aware reasoning.
In offline RL(Levine et al. 2020), where learning is based on sequences of states, actions, and rewards, Mamba was expected to outperform Transformers and other SSM-based decision models(Bar-David et al. 2023). However, integrating Mamba into decision models alone(Ota 2024) did not yield better performance compared to advanced variants of decision transformers.
We identified the underperformance as stemming from Mamba’s Selective Scan mechanism. Similar to attention, Selective Scan assigns weights to inputs, potentially excluding certain parts of the sequence from reasoning. Feature maps reveal that actions at specific steps are activated selectively, while others are neglected. This highlights the importance of combining information from adjacent steps to prevent complete omission before selection. Additionally, Mamba’s use of residual multiplication exacerbates information loss; values reduced near zero by Selective Scan remain negligible after residual multiplication, limiting their impact on outcomes.
Furthermore, as Mamba was originally developed for language modeling, it employs a single 1D convolution token mixer optimized for word embeddings. In offline RL, however, trajectories consist of states, actions, and returns-to-go (RTG), each possessing distinct characteristics. Using a single token mixer for all components hinders effective feature separation.
To address these challenges, we propose a novel approach termed Decision MetaMamba (DMM). Our approach incorporates a specialized token mixer to fuse information from adjacent steps, mitigating data loss. Additionally, we reposition the token mixer before the residual multiplication path and introduce a residual sum connection to counteract the loss of information caused by Selective Scan. Furthermore, by employing separate token mixers for states, actions, and RTG, we enhance feature extraction tailored to their unique properties.
Selective Scan has the advantage of selectively combining inputs to address the challenge of learning sub-optimal trajectories. Achieving an optimal path often requires “stitching” different trajectories, which previous studies have tackled using auxiliary networks(Wu, Wang, and Hamaya 2024)(Wang et al. 2024b) or search algorithms(Janner, Li, and Levine 2021). In contrast, DMM leverages Selective Scan to dynamically adjust trajectories by selectively focusing on specific input steps.
In experiments, we found that using a linear layer as a token mixer achieved comparable results to using 1D convolution, confirming that the critical factor is not the specific component but the ability to combine local information effectively.
We evaluated DMM on the D4RL(Fu et al. 2020) Mujoco, Antmaze, and Atari datasets. As shown in Figure 1, our model achieved competitive scores while using less than 10% of the parameters required by the state-of-the-art transformer model. Additionally, DMM outperformed existing models on most datasets in terms of overall scores.
By eliminating the need for positional encoding, Mamba reduces the risk of behavioral cloning associated with positional encoding in return-conditioned models and extends its applicability to scenarios where step information is unavailable. Since DMM inherits Mamba’s characteristics, it supports parallel scan(Nguyen 2007) for linear-time training and constant-time inference, similar to RNNs.
The main contributions of this study are as follows:
  • We demonstrated the importance of mixing input tokens within a window in offline RL decision models using Selective Scan SSMs. By incorporating a token mixer and residual network in the input layer, we achieved improved performance.
  • We designed the token mixer to account for different input modalities, enhancing performance while maintaining high efficiency with a reduced number of parameters.
  • By removing positional encoding from return-conditioned transformer decision models, we reduced the risk of behavioral cloning and extended the model’s applicability to datasets lacking proper time step annotations.
Related Work
Offline Reinforcement Learning
Offline reinforcement learning is a subfield of RL where the agent learns from a pre-collected dataset without additional environment interactions. This is crucial for applications where interactions are costly or risky. Challenges in offline RL include distribution shift and extrapolation error. Distributional shift is the mismatch between the data distribution in the offline dataset and the distribution the agent encounters when deploying the learned policy. Extrapolation error occurs when the policy queries state-action pairs that are not well represented in the dataset, which is often due to the lack of sufficient expert demonstrations in offline RL datasets.
Decision Model
Methodological advances to overcome hurdles in offline RL include Value-based behavioral regularization. Techniques like batch-constrained Q-learning (BCQ, (Fujimoto, Meger, and Precup 2019)), conservative Q-learning (CQL, (Kumar et al. 2020)), and implicit Q-Learning (IQL, (Kostrikov, Nair, and Levine 2021)) are designed to mitigate extrapolation errors by constraining the learned policy to stay close to the behavior policy that generated the dataset. Contrarily, Value-free approaches do not necessarily rely on Q functions. One such approach is Imitation learning, which aims to imitate the behavior policy by training on collected or desired trajectories filtered by heuristics or value functions. Trajectory Optimization, e.g. Multi-Game Decision Transformer (MGDT, (Lee et al. 2022)) and Trajectory Transformer (TT, (Janner, Li, and Levine 2021)), models joint state-action distribution over complete trajectories, reducing out-of-distribution action selection. To enable effective planning, this approach utilizes techniques such as beam search and reward estimates. Another widely known method is model-based approach. It uses learned models of the environment to generate synthetic data, which can augment the offline dataset and improve learning(Weissenbacher et al. 2022).
Information that can be obtained after an action is performed, such as returns-to-go, is called hindsight information, and it provides intuition on what should have been done differently. (Furuta, Matsuo, and Gu 2021) revealed that decision transformers learn policies through hindsight matching to ensure that the predicted statistical values for future states from the trajectory are satisfied.
However, trajectory stitching to find an optimal policy from suboptimal trajectories is challeging with hindsight matching alone. Consequently, various models utilizing an auxiliary network have been proposed. The Elastic Decision Transformer employed an auxiliary network to determine the optmial length of the input sequence for trajectory stitching. The Critic Guided Decision Transformer used a critic network to guide the policy to take actions consistent with the target returns. In contrast, the Decision Convformer enhanced performance by developing a token mixer to replace attention without using an additional network. This was inspired by the Metaformer, which demonstrated performance in vision tasks by replacing attention with a simple pooling layer. Beyond excelling in vision tasks, Metaformer has influenced numerous other studies by abstracting the token mixer to achieve versatile functionalities.
State Space Model
To advance language models further, new architectures capable of surpassing the performance of Transformers have been developed. Among these, SSM is a field originally utilized in control engineering for dynamic system mapping using state variables. In the domain of deep learning, SSM has been developed into models that learn by parameterizing system functions that transform discrete state variables(BOURDOIS 2023).
The input state x(t) is mapped to a latent space and then transformed back to the output state, which can be expressed by the following equation.
$\dot{h}(t)=Ah(t)+Bx(t)$ (1)
$y(t)=Ch(t)+Dx(t)$ (2)
Represented by linear differential equations, SSM includes three time-dependent variables.
$h(t)$ $\in\mathbb{C}^{n}\text{, n states variable}$
$x(t)$ $\in\mathbb{C}^{k}\text{, k states input}$
$y(t)$ $\in\mathbb{C}^{m}\text{, m outputs}$
This equation is represented by four learnable matrices.
$A$ $\in\mathbb{c}^{n\times n}\text{: state\,matrix, }B\in\mathbb{c}^{n\times k}\text{: control\,matrix,}$,
$C$ $\in\mathbb{c}^{m\times n}\text{: output\,matrix, }D\in\mathbb{c}^{m\times k}\text{: command\,matrix,}$,
$Dx(t)$ can simply be represented as a skip connection in the network, so it can be excluded during the stages of implementing SSM using the above equation and added later. Therefore, when deriving the equation that represents SSM as a convolution, $Dx(t)$ is excluded.
Finding the latent $h(t)$ from continuous signal input is analytically challenging. Moreover, as most problems involve discrete datasets, a discretized model is necessary for processing. Traditional methods for discretization include the trapezoidal method and the Newton method, while Mamba employs the zero-order hold (ZOH) technique. Discretization is one of the most crucial components in state-space models (SSMs) because it allows the transformation of differential equations handling continuous functions into recurrent models and convolution models. Mathematically, discretized learnable matrices $A$ and $B$ with zero-order hold(Gu and Dao 2023) can be represented as follows:
$\bar{A}=exp(\Delta A)$ (3)
$\bar{B}=(\Delta A)^{-1}(exp(\Delta A)-I)\cdot\Delta B$ (4)
With our SSM incorporating discretized input, it becomes possible to solve problems for specific timesteps. This approach can be divided into recurrent methods, such as RNNs, and parallelizable convolution methods. A model that appropriately combines these two approaches during training and evaluation is introduced in the Linear State-Space Layer (Gu et al. 2021).
The convolutive view involves diagonalizing the repeatedly operated matrix $A$ to create a convolution kernel through squaring.
Ultimately, the convolution kernel and representation can be expressed as follows.
$\text{kernel }K=(CB,CAB,CA^{2}B,..)\$ (5)
$y=K*x$ (6)
Here, the matrices $A$, $B$, and $C$, represented as the convolution kernel, remain identical regardless of the input sequence. Therefore, we obtain a static representation model that is not content-aware. Among these, matrix $A$ plays the most crucial role as it compresses information from the previous state and conveys it to the new state. To efficiently and extensively store information from the previous states, $A$ is constructed using Hippo (High-order Polynomial Projection Operator). The imposition of Hippo as a structural condition on matrix $A$ results in the Structured SSM, known as S4 (Gu, Goel, and Ré 2021).
Figure 2: Offline RL learns from different trajectories. To learn the optimal policy, trajectory stitching must be performed by following the path of trajectory A until step t and then switching to the path learned from trajectory B at step t+1. In models that only attend to the current step, the accuracy of actions can be degraded due to the influence of multiple overlapping trajectories. In contrast, considering the influence of distant steps appropriately can facilitate the prediction of optimal actions by combining with the return condition.
Mamba
Mamba(Gu and Dao 2023) developed its model based on the structural foundation of S4 through two significant modifications. Previously, the linear time-invariant matrices $A$, $B$, and $C$ were the same for all input tokens. As a result, it was unable to solve the selective copying problem, lacked content-aware reasoning, and consequently, could not achieve significant performance improvements in language tasks and performed poorly on the induction heads problem.
To address these issues, Mamba allowed matrices $B$ and $C$, as well as the step size $\Delta$, to vary based on the input. This enabled the model to possess content awareness. However, since $B$ and $C$ are now dynamically selected, it is no longer possible to use convolution parallelism with fixed kernels. These states can be easily computed through sequential operations using a for loop. However, parallelization presents a different challenge. What initially seemed impossible to parallelize was achieved by Mamba using a parallel scan algorithm (Harris, Sengupta, and Owens 2007).
Figure 3: The network architecture of Decision MetaMamba (DMM) modifies the input layer of the Vanilla Mamba structure. By utilizing a multi-modal token mixer for sequence transformation at the forefront of the input layer, the model is designated as a Single DMM. If an additional multi-modal token mixer is added after the input projection, it becomes a Double DMM model. In the Vanilla Mamba architecture, a 1D convolution token mixer is used after the input projection, whereas in the Single DMM, the token mixer is positioned before the input projection to ensures that all inputs of the SSM receive an integrated input that considers short sequences.. ’LL’ denotes Linear Layer.
Additionally, Mamba improved its speed by efficiently utilizing the hardware architecture of GPUs. By employing kernel fusion and recomputation of immediate states during the backward pass, it minimized data transfer between SRAM and DRAM (HBM).
Method
In an ideal scenario, an environment satisfying the Markov property would allow for the prediction of optimal paths using only the current state within a sequence. However, in practice, utilizing a shorter sequence results in a reduction in the accuracy of action prediction during both training and inference. Introducing a longer sequence input, as illustrated in Figure 4, makes it challenging to select an action that deviates from the current path while achieving a higher return. To address this, the Elastic Decision Transformer was devised to incorporate an auxiliary network for adapting the length of the input sequence during inference. Mamba’s selective scan is capable of selectively integrating the information of the input sequence during inference. This results in an effect akin to inputting shorter steps by only selecting a portion of the input sequence. However, this process could potentially lead to the omission of critical information. The DMM was developed to adopt the token mixer within Mamba’s input layer and connect a residual sum network to the SSM’s output, thereby maintaining significant local relationships between adjacent steps while simultaneously considering the input selectively to infer the next action. A residual sum connection was incorporated within the Mamba block to alleviate learning difficulties caused by near-zero values, which could not be rectified through residual multiplication in the Mamba block. The token mixer was developed in two variations, utilizing 1D convolution and linear layers, to examine its impact.
Model Architecture
As illustrated in Figure 3, the original Mamba is divided into three main components. The first component is the input layer, which transforms the input state using a combination of linear layers, 1D-convolution layers, and activation functions. The second component is the SSM layer. As detailed earlier, the selective scan SSM performs content-aware reasoning while enabling fast, parallel operations through a hardware-aware algorithm. The final component is the output layer, which linearly transforms the output of the SSM.
Among the layers in our architecture, we have incorporated a token mixer into the input layer. Following the token mixing process, layer normalization is applied, and the output is passed to the SSM. A residual network is added after layer normalization and is merged with the output of the SSM layer. The offline reinforcement learning (RL) dataset consists of three independent modalities: returns-to-go, state, and action. These scalar values are transformed into embedded vectors using a linear layer. Depending on the configuration, these values are mapped to vectors of a predetermined size, which can be adjusted based on the dataset. In our experiments with DMM, we used 64-dimensional vectors. At this embedding stage, we employ different linear networks to process the data. Since the dataset contains different modalities, we also utilize distinct token mixers to preserve the characteristics of each modality after token mixing. We developed and evaluated the token mixer in two forms: a multi-modal 1D convolutional layer and a multi-modal linear layer. The rationale for using two forms is to demonstrate that performance is not reliant on a specific component. Instead, we show that performing token mixing prior to selective scanning is a generalizable method for enhancing performance. We also explored integrating a new structure into Mamba’s input layer, aiming to establish a method that can be applied to similar approaches when utilizing Mamba in other applications. Although the input layer is not the core structure of Mamba, we suggest that abstracting and incorporating various configurations could contribute to performance improvements.
Double Token Mixer
The original Mamba structure includes a 1D convolutional token mixer positioned before the SSM. In DMM, we experimented with two modifications: either retaining or removing this token mixer. When both a token mixer and a residual network were added to the very beginning of the input layer, while retaining the original token mixer before the SSM, this configuration was termed the ”Double Token Mixer” model. The additional token mixer employed the same multi-modal 1D convolution or linear layer as the first token mixer. By introducing this additional token mixer, the token mixing process is performed once again before passing inputs into the SSM. Additionally, the output of this new token mixer is multiplied with the output of the first token mixer, creating a synergistic interaction between the two layers.
Figure 4: Multi-modal 1D convolution and linear layer are employed to extract hidden states using different layers for Rtg, state, and action embeddings. 1D-convolution connects different modes in the hidden dimension, while the linear layer connects them in the sequence dimension.
Input Token Mixers
In the input layer, both token mixers—1D convolution and linear layer—perform sequence transformation while considering causality. Consequently, they do not include information beyond the current time step in the filter. To achieve this, zero padding of $window\_size-1$ is added to the beginning of the sequence vectors.
Multi-Modal 1D-Convolution
The convolution module aims to integrate neighboring embeddings within the window size. It applies filters across channels of the hidden dimension using a 1D convolution kernel. As illustrated in Figure 4a, to account for multi-modality, three different convolution kernels are used. This results in three distinct hidden states, generated by each kernel in the order of Rtg, state, and action. These hidden states are then combined to form a new hidden state tensor. The output tensor size matches the input tensor size.
1D convolution cannot account for positional information between inputs. Although methods such as (Sabour, Frosst, and Hinton 2017) have been developed to address this issue, they are too computationally intensive for decision models. With a filter length of 6, only two time steps $(rtg_{t-1},state_{t-1},action_{t-1},rtg_{t},state_{t},action_{t})$are considered. Although information outside the convolution window can be integrated through layer stacking, it is challenging for the entire information to be encompassed within the final step hidden state.
Multi-Modal Linear
As can be observed in Figure 4b, Zero-padded sequence vectors are grouped into embedding sets of window size. To perform computation in a single linear transformation without iteration, each window size vector within the set is flattened into a 1-dimensional vector. These are then transformed back into vectors with the original hidden state dimension via a linear layer. While the convolution module applies filters to the hidden state dimension, the linear module concatenates consecutive sequence vectors of window size in the time domain in a flattened manner. The resulting vectors, transformed back to the hidden state dimension, form a sequence, yielding an output tensor that matches the input tensor size. The linear layer also uses three different modules to distinguish between each modality.
Train and Evaluation
Train
In the train trajectory dataset, each sequence of length, which includes Rtg, state, and action, is embedded into vectors of the same size through an input embedding layer. This process addresses the issue arising from the differing sizes of these components. Traditional transformer architectures typically add positional encoding to the embeddings. However, DMM ensures input causality through padding and processes the inputs using a sequence model, thereby obviating the need for positional encoding. By eliminating additional information dependent on the trajectory, the likelihood of behavioral cloning is further reduced. The output from passing through $n$ MetaMamba blocks is transformed into an action-dimension vector, resulting in the predicted action. The loss is then computed between the predicted actions and the true actions across all sequence lengths. As described, the loss is calculated for all predicted actions within a sequence, but during evaluation, only the final action in the sequence length is used.
$\mathcal{L}_{DMM}=\mathbb{E}_{\tau\in\mathcal{D}}\left[\frac{1}{K}\sum\limits_{t=1}^{K}(a_{t}-\hat{a}_{t})\right].$
While a longer sequence length during training can provide more guidance, it may inversely affect the accuracy of the final action. In a situation where there is a shortage of necessary guides for learning, DMM was able to enhance the accuracy of the final action by considering both short-range and long-range information within the sequence. Therefore, DMM utilized shorter sequence lengths in all environments compared to existing models.
dataset type TD3+BC CQL DT EDT DC DS4 DM DMM-L DMM-C
hopper md 59.3 58.5 68.4 70.4 89.7 54.2 86.2 96.2 94.2
hopper mr 60.9 95.0 82.5 89.4 84.8 49.6 81.7 95.0 91.2
hopper me 98.0 105.4 108.7 104.7 110.1 110.8 111.0 110.2 109.9
walker2d md 83.7 72.5 76.6 75.8 79.8 78.0 77.6 81.3 81.6
walker2d mr 81.8 77.2 65.9 73.3 78.2 69.0 72.1 74.1 73.9
walker2d me 110.1 108.8 109.0 107.8 109.2 105.7 108.3 110.2 109.5
halfcheetah md 48.3 44.0 42.8 43.0 43.0 42.5 42.8 43.0 43.3
halfcheetah mr 44.6 45.5 39.1 37.8 41.1 15.2 39.8 41.1 41.3
halfcheetah me 90.7 91.6 85.5 89.7 93.1 92.7 90.6 92.6 92.5
locomotion mean 75.3 77.6 75.4 76.9 81.0 68.6 78.9 82.6 81.9
antmaze um 78.6 74.0 69.2 67.8 86.0 63.4 68.4 88.0 85.0
antmaze ud 71.4 84.0 61.0 58.3 81.0 64.6 62.1 80.0 78.0
antmaze mean 75.0 79.0 65.1 63.1 83.5 64.0 65.3 84.0 81.5
Table 1: We summarize the experimental results using the expert normalized return score according to (Fu et al. 2020). The experiments recorded the highest value from results using 5 random seeds. The results highlighted in bold indicate the highest or comparable scores in each benchmark. DMM uses two variations of input projection layers, with ’L’ representing multi-modal linear and ’C’ representing multi-modal convolution layers.
Evaluation
In the evaluation stage, DMM interacts with the environment in real-time to infer the action token. Therefore, it is necessary to set the initial Rtg. Consistent with (Chen et al. 2021)., we set the initial target return to be higher than the actual return of the trajectory.
Experiments
We examined whether DMM can demonstrate improved performance by considering both proximate and distance sequences. This performance evaluation utilized pre-collected datasets from the D4RL (Fu et al. 2020) MuJoCo, AntMaze, and Atari (Mnih et al. 2013) domains. Additionally, we employed two modules that replace the input layer of SSM to ascertain if DMM can enhance offline RL performance compared to models using traditional SSMs. Since SSMs were originally developed to replace self-attention in LLMs, they have previously been used to substitute the token mixer in decision transformers. Contrarily, through the selective scan SSM and an abstracted input layer structure, DMM suggests a future direction for decision models utilizing SSM.
Dataset DT DC DMM-DL DMM-DC
hopper-md 68.4 85.6 98.6 96.5
hopper-mr 78.5 89.5 95.9 92.1
hopper-me 101.3 101.5 110 109.5
walker2d-md 81.1 82.6 81.4 82.2
walker2d-mr 49.2 69.3 75.1 78.9
walker2d-me 104.8 104.7 110.4 109.8
halfcheetah-md 42.9 42.8 43.2 43.3
halfcheetah-mr 37.5 38.7 41.1 41.5
halfcheetah-me 82.8 86.3 92.6 92.2
antmaze-um 64.4 81.0 80.0 82.0
antmaze-ud 58.0 68.0 74.0 77.0
Table 2: Offline performance results of DMM and baselines in the Mujoco and Antmaze domains. In this experiment, all models employed an input embedding of 64 dimensions. The results indicate that DMM achieved improved performance compared to existing models even under conditions with fewer parameters.
D4RL Mujoco and Antmaze Environments
We measured the model’s performance using two offline RL datasets provided by D4RL: Mujoco and Antmaze. Mujoco is an environment for controlling moving objects, utilizing more state and control variables compared to classic control from (Brockman et al. 2016). The objects with mobility receive rewards based on duration and state, thus receiving continuous rewards across the entire sequence. In contrast, Antmaze, by default, is set to a sparse reward setting, where rewards are only received upon reaching the final target position.
Game DT DC DM DMM-DL DMM-DC
Breakout 284.5 334.3 256.9 202.4 362.2
Pong 106.8 109.6 104.0 104.0 106.8
Qbert 38.1 64.4 47.6 59.1 63.6
Seaquest 2.8 2.6 2.2 2.6 2.7
Table 3: Offline performance results of DMM and baselines in the Atari datasets. The experiment yielded the highest result using three fixed values as random seeds.
Results
The selection of baseline models for comparison was divided based on the structure adopted by each model, comprising a total of seven models: Value-based RL models TD3+BD(Fujimoto and Gu 2021), CQL(Kumar et al. 2020), models with Transformer architecture DT(Chen et al. 2021), EDT(Wu, Wang, and Hamaya 2024), DC(Kim et al. 2024), and models utilizing SSM DS4(Bar-David et al. 2023), DM(Ota 2024). The results are a combination of experimental results conducted in this study and cited results. For detailed information, please refer to Appendix C.
Experimental results predominantly indicated enhanced or comparable performance across environments. Above all, it is significant in that it demonstrates a higher performance improvement compared to previously proposed offline RL models utilizing SSMs. (Bar-David et al. 2023) lacked the selective scanning feature, which prevented it from utilizing inputs selectively in inference. Additionally, (Ota 2024) did not have an input layer capable of effectively mixing tokens.
Moreover, compared to models incorporating a Transformer architecture, a marked performance improvement was observed specifically within the hopper. This highlighted the potential of models designed to integrate both proximate and distant sequences, which can be achieved through the effective combination of SSM and multi-modal token mixers
The performance varied when the multi-modal layer was implemented as a 1D convolution layer versus a linear layer, depending on the environment. This indicates that, similar to Metaformer(Yu et al. 2022), adjusting the input layer according to the domain can lead to performance improvements. Based on this, various input layers can be developed to enhance performance. However, in almost all benchmark tests, using a multi-modal layer resulted in better outcomes compared to using a layer that mixes embeddings without distinguishing between input modalities. This underscores the importance of considering multi-modality in the process of interpreting sequences and selecting actions, and applying this to the modification of abstracted input layers.
Ablation Study
To validate the model architecture, we conducted new ablation studies. The first study tested the model without the input layer, isolating the impact of the added layer by retaining the second 1D convolution in the Double input sequence transformation model. The highest normalized test score was recorded after five random seed trainings. Performance degraded in all benchmarks when the input token mixer was excluded.
Dataset DMM-DC DMM-DL DMM-DC w/o input layer DMM-DL w/o input layer
hopper-md 96.5 98.6 88.3 89.5
walker2d-md 82.2 81.4 77.3 78.6
antmaze-um 82.0 80.0 68.0 66.0
Table 4: Performance Comparison with or without input layer
The second study excluded the token mixer for multi-modality (state, action, returns-to-go). With a modest reduction in parameters, performance declined, though not as much as when the input token mixer layer was removed. We will include these results in the appendix to support our claims.
Dataset DMM-DC w/o multi-modal token mixer DMM-DL w/o multi-modal token mixer
hopper-md 93.7 94.4
walker2d-md 80.9 81.2
Table 5: Performance without multi-modal token mixer
Conclusion
We propose Decision Metamamba, a sequential decision-making model that employs the selective scan State-Space Model with a modified input layer. The input sequence in offline reinforcement learning includes data in different formats: state, returns-to-go, and action. To handle this, we utilized two distinct token mixer networks, collectively referred to as the Multi-Modal Token Mixer. These two token mixers are, respectively, token mixers that utilize a 1D convolution layer operating in the hidden state dimension and a linear layer for integrating tokens along the sequence dimension. Each network demonstrated strong performance in specific experimental domains.
To appropriately consider information from both short sequences close to the current step and distant steps, we employed the selective scan SSM. Unlike previous sequence models, the Selective Scan SSM facilitates rapid learning through parallel computation while utilizing a relatively small number of parameters, as some layers perform convolutional operations. This combination allows the model to achieve scores comparable to or exceeding those of existing models, despite having a lower number of parameters. Moreover, leveraging the recurrent properties of the transition matrix allows for inference speeds up to five times faster than traditional transformer models. Additionally, since the output of the token mixer sequentially serves as the input to the SSM for action inference, the model does not require time-step encoding, contributing to its enhanced generalizability.
The ongoing advancements from the initial Structured SSM to Gated SSM and Recurrent SSM (Patro and Agneeswaran 2024a) suggest that this study can pave the way for improved performance in offline RL decision models through the use of advanced SSMs.
Appendix
A. Self-Attention and SSM
We have verified the impact of MMTM on performance through previous experiments. Additionally, to investigate whether SSM can improve performance when used with other types of token mixers, we tested a model using self-attention from Decision Transformer as the token mixer and Mamba’s SSM as the channel mixer, replacing the conventional MLP layer. The results, as shown in Table 4, demonstrated a remarkable improvement in performance.
Both Self-attention and selective scan SSM operate by selectively combining inputs to generate outputs, but they employ different methods. This combination of their strengths led to enhanced results. However, we obtained lower performance compared to models utilizing 1D convolution or linear layers. This observation suggests that the concept of the DMM, which first integrates tokens in close proximity and then performs selective integration with information from tokens at distant steps, can provide higher performance compared to models that only employ two types of selective combinations.
Self-Attention Self-Attention + SSM
hopper-md 68.4 83.6
hopper-mr 82.5 90.3
Table 6: Performance Impact on Self-attention token mixer model with SSM as a Channel Mixer
B. Parameter Efficiency
As shown in Figure 2, the DMM requires significantly fewer parameters to train a model of equivalent performance compared to traditional models using the transformer architecture. This is attributed to the model’s characteristics that utilize the Selective Scan SSM, which employs convolutional filters, and its ability to maintain performance with smaller input embeddings. Table 7 presents the total number of parameters and the parameter-to-score ratio for all the models we experimented with. When using MMTM entirely as 1D-Convolution, the overall model size could be reduced to approximately 6% while achieving similar performance. This demonstrates the achievement of developing a resource-efficient model. Even when employing a linear layer for the input layer, the model still exhibited efficiency in the parameter-to-score ratio.
Dataset item DT DC DMM-C DMM-DC DMM-L DMM-DL
hopper-md $n$ of params 727695 1858307 358403 110467 1235075 1204291
hopper-md $n$ of params / score 10638.8 20716.9 3804.7 1144.7 12838.6 12213.9
walker2d-md $n$ of params 730008 537478 32390 36422 85670 306854
walker2d-md $n$ of params / score 9530.1 6735.3 396.9 443.1 1053.8 3769.7
antmaze-um $n$ of params 730008 539528 361992 112264 1238664 1206088
antmaze-um $n$ of params / score 10549.2 6273.6 4258.7 1403.3 14075.7 14708.4
Table 7: parameter efficiency
C. Implementation Details
All experiments were conducted using an Nvidia RTX 3060 12GB GPU on an Ubuntu OS.
In the experiments, to limit the effect of hyperparameters, most parameters were used identically to those in the (Kim et al. 2024) except the embedding dimension.
The results of the value-based model in the Mujoco and Antmaze domains are cited from Decision Convformer, while the Antmaze results for EDT are cited from (Zhuang et al. 2024). The remaining results are recorded from our own experiments.
Name Value
Number of layers 3
Batch size 64
Context length $K$ 8
Embedding dimension 64 (32 for Walker2D)
Dropout 0.1
Nonlinearity function SiLU
Grad norm clip 0.25
Weight decay $10^{-4}$
Learning rate decay Linear warmup
Total number of updates $10^{5}$
Table 8: Common hyperparameters of DMM on training of Mujoco and Antmaze
As confirmed by the experimental results in the Table 7, when the input embedding dimension was reduced, the DMM model exhibited robust results without performance degradation, unlike other models. Thus, the adjusted input dimension was set as the default.
DMM-DL 128 DMM-DL 64
hopper-md 97.1 96.2
walker2d-md 79.7 81.3
halfcheetah-md 43.1 43.0
Table 9: Performance According to Input Embedding Dimension
In the Atari domain, the setup was generally consistent with that of the Decision Convformer except embedding dimension. (Kim et al. 2024) employed embeddings of 256 dimensions in the hopper to enhance performance, which resulted in a substantial difference in the number of parameters used, as illustrated in Figure 2, when compared to the DMM.
Name Value
Number of layers 6
Embedding dimension 128
Batch size 64
Return-to-go conditioning 90 Breakout ($\approx$ 1 $\times$ max in dataset)
2500 Qbert ($\approx$ 5 $\times$ max in dataset)
20 Pong ($\approx$ 1 $\times$ max in dataset)
1450 Seaquest ($\approx$ 5 $\times$ max in dataset)
520 Asterix ($\approx$ 5 $\times$ max in dataset)
950 Frostbite ($\approx$ 5 $\times$ max in dataset)
780 Assault ($\approx$ 5 $\times$ max in dataset)
2750 Gopher ($\approx$ 5 $\times$ max in dataset)
Encoder channels 32, 64, 64
Encoder filter sizes $8\times 8$, $4\times 4$, $3\times 3$
Encoder strides 4, 2, 1
Max epochs 10
Dropout 0.1
Learning rate $6\times 10^{-4}$
Adam betas (0.9, 0.95)
Grad norm clip 1.0
Weight decay 0.1
Learning rate decay Linear warmup and cosine decay
Warmup tokens 512 * 20
Final tokens 2 * 500000 * $K$
Context length 20
Table 10: Common hyperparameters of DMM on training of Atari
D. The Effect of Activation Functions RELU, GELU, and SiLU.
We confirmed performance improvements by changing the activation function to GELU during experiments with the Decision Convformer on the AntMaze task, which provides sparse rewards. In the Transformer architecture, the ReLU activation function is typically used following the MLP, which acts as a channel mixer. The modification involved replacing this with GELU. Conversely, in the Mamba architecture, the SiLU activation function is used in the input layer preceding the selective scan SSM.
As observed in the Table 9, changing the activation function from SiLU to GELU in the DMM did not result in any performance change. The reason for this can be found in Figure 5, which shows that SiLU and GELU exhibit similar functional behavior. The functional difference between ReLU and GELU is most pronounced near zero, where the majority of values lie, potentially impacting performance. However, since SiLU does not exhibit significant differences in function values compared to GELU around zero, the impact on overall model performance is likely to be minimal.
Figure 5: Graphs of ReLU, GELU, SiLU activation functions
DMM-GELU DMM-SiLU
antmaze-um 82.4 85.0
antmaze-ud 78.0 81.0
Table 11: Performance According to Input Embedding Dimension
E. Selective Scan Feature Activation
We examined whether the hidden state at a specific step could approach zero by analyzing the output feature map of the selective scan. As shown in Figure 6, the values at certain steps were predominantly distributed near zero. Due to the structure of Mamba, where the output of the selective scan SSM is multiplied by the residual, causing the values to become zeros, it becomes challenging for signals to propagate effectively. DMM addresses this issue by incorporating a residual sum mechanism. Furthermore, we observed that the actions at specific steps were activated in the feature map, confirming that information from distal steps is integrated during the final inference process.
Figure 6: A feature map from a selective scan SSM